OpenStudy (anonymous):

Find all solutions of the equation: cos x/2 = (sq rt 2)/2

OpenStudy (anonymous):

second primary lesson :) cos(x/2) = sqrt(2)/2 The first thing you can do is take the inverse cosine (arccos) of the other side, leaving you with: x/2 = arccos(sqrt(2)/2) Then you multply both sides by 2, giving you: x = 2*arccos(sqrt(2)/2) Now, the inverse cosine will return a value from 0 to Pi, so x will be from 2*0 to 2*Pi, or 0 to 2Pi. So, I think you just have to plug that into your calculator to find x.

OpenStudy (anonymous):

resolves?